# How to Collaborate: Towards Maximizing the Generalization Performance in Cross-Silo Federated Learning

### v1

@@ 5.1 Discussions
In some scenarios, the result of client partition in HCCT can degenerate to the following collaboration patterns:
@@
Connection to personalized FL. The proposed HCCT scheme is orthogonal to the personalization techniques in FL [9, 10, 22, 23, 24]. Specifically, each group in HCCT can be viewed as “a micro FL system” that contains a set of clients. Thus, we can employ any personalization technique to enhance the clients’ performance in a group. For example, the clients may only collaborate to train the feature extractor of a model and keep the classifier as a personalization layer locally. In Section 7.3, we demonstrate the compatibility of HCCT with personalization techniques via simulations.
@@ 5.2 Efficient Estimation of Client Similarity
Computing client utility requires repeatedly evaluating client similarity $s\left(i,G_{i}\right)$. Despite that this operation is performed on the central server, it still introduces high computation complexity of $\mathcal{O}(M)$ due to the vector multiplication between two $M$-dimensional gradients. To reduce such complexity, we propose an efficient implementation by finding the most important layer accounting for client similarity. After the first training iteration, the server collects the gradients from all clients. The gradient of client $i$ can be expressed in a layer-wise manner as follows:
@@
This efficient implementation reduces the computation complexity from $\mathcal{O}(M)$ to $\mathcal{O}(dim(\mathbf{g}_{i}^{t}[l^{*}]))$.
@@ 6 Convergence Analysis
In this section, we will analyze the proposed HCCT scheme to prove its convergence and observe the effect of client similarity. To begin with, we make some commonly adopted assumptions on the local loss functions [41, 42, 43, 44]. It is worth noting that we do not assume any convexity of loss functions in this section. In other words, the following results hold for general non-convex loss functions. For simplicity, we denote the local empirical loss function of client $i$ by:
@@
$\mathbb{E}\left[\sum_{(\mathbf{x},y)\in\mathcal{B}}\nabla_{\mathbf{w}}l(h((\mathbf{x};\mathbf{w}),y))\right]=\nabla J_{i}(\mathbf{w}),\forall i\in\mathcal{N}.$ (23)
@@
$\mathbb{E}\left[\left\|\sum_{(\mathbf{x},y)\in\mathcal{B}}\nabla_{\mathbf{w}}l(h((\mathbf{x};\mathbf{w}),y))-\nabla J_{i}(\mathbf{w})\right\|_{2}^{2}\right]\leq\sigma^{2},\forall i\in\mathcal{N}.$ (24)
The data distribution on clients varies among each other, as explained in Section 3.1, which leads to different optimization objectives and diverse gradients. To show such data heterogeneity, previous works [45, 46] made assumptions on the difference between local and global training objectives. Different from these works, we aim to investigate the collaboration among clients and thus make a more precise assumption on the gradient similarity among clients as follows.
@@
We are now to derive the convergence of the HCCT in the following theorem. Previous works in FL [47, 5, 48] upper bound the expected squared gradient norm of the global loss function. Compared with these works, the clients in cross-silo FL have local training objectives and are not concerned about the global objective. Therefore, a better option is to verify the convergence of the algorithm using the sum of local gradients $\sum_{i\in\mathcal{N}}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})\|^{2}]$ [49].
@@
Let $\xi_{T}=\sum_{t=0}^{T-1}\eta^{t}$. With Assumptions 1-2, and 3, if the learning rates satisfy $\eta^{t}<\frac{1}{L},\forall t\in[T]$, we have:
@@
$+\frac{1}{\xi_{T}}\sum_{t=0}^{T-1}\frac{(\eta^{t})^{3/2}}{2}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$
$+\frac{1}{\xi_{T}}\sum_{t=0}^{T-1}\frac{L^{2}t}{2\sqrt{\eta^{t}}}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{i\in\mathcal{N}}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}$
$+\frac{1}{\xi_{T}}\sum_{t=0}^{T-1}L(\eta^{t})^{2}N\sigma^{2}+\sum_{t=0}^{T-1}\frac{L^{2}Nt}{\sqrt{\eta^{t}}}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2}.$ (26)
@@
The key step of the proof is to upper bound the divergence between any two local gradients, i.e., $\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$. Using $L$-smoothness in Assumption 1, we show that:
@@
Then we observe that, despite evolving from different groups, $\mathbf{w}_{j}^{t}$ and $\mathbf{w}_{i}^{t}$ have the same initialization $\mathbf{w}^{0}$. Therefore, we can upper bound their difference by comparing the accumulated gradients in previous $t-1$ iterations. The detailed proof is deferred to Appendix C. ∎
@@
When the learning rates satisfy $\lim_{T\rightarrow\infty}\sum_{t=0}^{T-1}\eta^{t}=0$ and $\lim_{T\rightarrow\infty}\sum_{t=0}^{T-1}(\eta^{t})^{3/2}>0$, the right-hand side (RHS) of (26) converges to zero as $T\rightarrow\infty$, i.e., the local models output by HCCT converge to the stationary points of the local loss functions.
From Theorem 2, we observe that the convergence of HCCT is hindered by the gradient dissimilarity among clients. Specifically, as more heterogeneous clients are grouped, i.e., $\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$ is larger, the convergence speed will be slowed down. This verifies our intuition that clients with similar gradients should be divided into the same collaboration group. In the next section, we will show the empirical benefits of the proposed HCCT scheme via extensive simulations.
@@ 7.1 Setup
  • CIFAR-10 [34]: In this training task, $N=10$ clients have heterogeneous and imbalanced data samples. We divide the training dataset into 100 shards, each of which contains one random class of data samples. Then, we randomly allocate a certain number of shards to each client using a half-normal distribution $\mathcal{H}(1)$.
@@
  • FedFA [^2] This work [25] adopts SGD with momentum as local training algorithm, but for fair comparisons we use mini-batch SGD in all methods. [25]: All clients train a global model similarly to the global training scheme. The key difference is that their model updates are weighted according to the training accuracy ${A}_{i}^{t}$ on local data, i.e., $\hat{\mathbf{g}}_{1}^{t}=\sum_{i\in\mathcal{N}}\frac{-\log_{2}{A}_{i}^{t}}{\sum_{j\in\mathcal{N}}-\log_{2}{A}_{j}^{t}}\mathbf{g}_{i}^{t}$, which ensures equitable accuracy distribution among clients.
@@
To demonstrate the compatibility of HCCT, we also show the simulation results for two variants: HCCT-E (an efficient implementation as in Section 5.2) and HCCT-P (with personalization layers [23]).
@@ 7.2 Performance Comparison
In the FMNIST training task, HCCT and global training have similar test errors, and they outperform other baselines. Given the IID data at clients, a global collaboration incorporates the maximum training data samples and significantly improves the generalization performance of the model. However, clients in the CIFAR-10 task have heterogeneous data, and thus the global training scheme suffers from such non-IID issue. In this case, it is preferable for clients to perform independent training, and HCCT also exhibits comparable test error to independent training. These results evidence that HCCT can identify the most effective client collaboration pattern, including independent training and global training, to accommodate various scenarios.
To further compare different training schemes, we summarize the average, standard deviation, minimal, and maximal values of test errors at clients in the Digit training task in Table V. We see that in addition to achieving the lowest average test error, HCCT also notably reduces the test error of the worst-performing client, which verifies that clients with fewer training data samples can benefit from effective collaboration. Meanwhile, the generalization performance of those clients with sufficient training data is also preserved, as the proposed utility function excludes clients with dissimilar data from its collaboration group. Besides, it is worth noting that HCCT achieves a lower standard deviation than the baseline schemes, which shows that HCCT implicitly enhances fairness among clients, leading to similar generalization performance for all clients.
@@ 7.4 Effect of System Parameters
Effect of $\alpha$. In Fig. 4, we investigate the effect of $\alpha$ using the CIFAR-10 dataset. A larger value of $\alpha$ emphasizes the importance of data volume, while a small value means that clients tend to collaborate with someone with similar data distribution. In this study, we split the assigned data at each client into the training data and test data according to different ratios, where a larger ratio means that a client has more training data samples but less test data. We observe from the results that as the ratio becomes larger, more training data speed up the training process and enhance the generalization performance in all cases. Moreover, if clients have fewer training data samples (i.e., ratio is $0.05$ or $0.1$), the optimal value of $\alpha$ is $100$. It means that clients have limited local data and tend to collaborate with others to benefit from their training data. In contrast, when clients have many local data samples (i.e., ratio is $0.2$ or $0.5$), they will emphasize more on the data distribution. Thus, by using a smaller $\alpha$ clients can avoid collaborating with other clients whose data distribution diverges significantly. These results provide a guideline for the selection of the value of $\alpha$ in the utility function. In addition, it is straightforward to extend and apply personalized values of $\alpha$ for different clients.
@@
Effect of the number of groups. Next, to study how the number of groups affects the generalization performance, we set the number of clients as $N=20$, manually fix the number of groups as $K$, and evaluate the test error in the Digit training task. From the results in Fig. 5, we observe that in this scenario the optimal number of groups to achieve the minimal test error is $10$. Notably, the maximal utility is also achieved when clients are divided into $10$ groups. A plausible explanation is that each group consists of clients sharing the same dataset. This implies that the utility function can effectively reflect the client’s requirement of improving the generalization performance, such that the minimal test error can be achieved when the utility is maximized.
@@ 8 Conclusions
In this work, we tackled the challenge of optimizing the client collaboration pattern in order to maximize the generalization performance in cross-silo FL. We derived the generalization bound for clients in various collaboration cases and then formulated the client utility maximization problem. To efficiently solve this problem, we proposed HCCT, a hierarchical clustering-based collaborative training scheme, in which clients are partitioned into different non-overlapping groups without the need to initially decide the number of groups. We also proved the convergence of HCCT for general loss functions. The effectiveness of HCCT is further verified via extensive simulations in different scenarios and datasets. For future work, it is worth exploring personalization techniques in HCCT to further improve the generation performance.

### v2

@@ 5.1 Discussions
In some scenarios, the result of client partition in HCCT can degenerate into the following collaboration patterns:
@@
Connection to personalized FL. The proposed HCCT scheme is orthogonal to the personalization techniques in FL [9, 10, 22, 23, 24]. Specifically, each group in HCCT can be viewed as “a micro FL system” that contains a set of clients. Thus, we can employ any personalization technique to enhance the clients’ performance in a group. For example, the clients may only collaborate to train the feature extractor of a model and keep the classifier as a personalization layer locally. This variant, named HCCT-P is advantageous when clients’ data are more diverse, as incorporating personalized layers can enhance client-specific model accuracy. As such, HCCT-P is tailored to scenarios characterized by a high degree of non-IID data among clients. In Section 7.3, we demonstrate the compatibility of HCCT with personalization techniques via simulations.
Computational complexity. The proposed HCCT scheme incurs only the computation cost on the server without bringing any additional overhead to clients. Specifically, the server needs to compute client similarity to evaluate the client’s utility. The computational complexity of client similarity is related to the number of clients and the number of parameters in the gradient, as detailed in the following. On one hand, the first step of the client partition function requires computing gradient similarity between two clients, which is repeated for $\frac{1}{2}N(N-1)$ times. In the best case, when there is no benefit for merging any two groups, the computational cost is $\mathcal{O}(N(N-1))$. Otherwise, in each step $2\leq l\leq N$, there is a newly generated group, which consists of at most $l+1$ clients, and other $N-l-1$ groups. Since we iterate over other groups and compute the benefits if it is merged with this new group, this process results in a computational cost of $\mathcal{O}(N(l+2)-(l+1)^{2})$. In the worst case, where we continuously put clients into one single group until it becomes a global training, the computational complexity is $\mathcal{O}(\frac{1}{6}(N^{3}-19N+6))$. It is worth noting that the number of clients $N$ in cross-silo FL is rather small [4], leading to an acceptable cost on the server. On the other hand, since the gradient similarity is computed as the vector multiplication between two $M$-dimensional gradients, it introduces a computational complexity of $\mathcal{O}(M)$.
@@ 5.2 Efficient Estimation of Client Similarity
Computing client utility requires repeatedly evaluating client similarity $s\left(i,G_{i}\right)$. Despite that this operation is performed on the central server, it still introduces high computational complexity of $\mathcal{O}(M)$ due to the vector multiplication between two $M$-dimensional gradients. To reduce such complexity, we propose an efficient implementation by finding the most important layer accounting for client similarity. After the first training iteration, the server collects the gradients from all clients. The gradient of client $i$ can be expressed in a layer-wise manner as follows:
@@
This efficient implementation reduces the computational complexity from $\mathcal{O}(M)$ to $\mathcal{O}(dim(\mathbf{g}_{i}^{t}[l^{*}]))$.
To summarize, HCCT-E is an efficient implementation of HCCT, designed to reduce the computational complexity in the client partitioning process. This is particularly advantageous in environments where clients have similar data distributions and dataset sizes, as a coarse partition of clients is sufficient.
@@ 6 Convergence Analysis
In this section, we analyze the proposed HCCT scheme to prove its convergence and observe the effect of client similarity. To begin with, we make some commonly adopted assumptions on the local loss functions [43, 44, 45, 46]. It is worth noting that we do not assume any convexity of loss functions in this section. In other words, the following results hold for general non-convex loss functions. For simplicity, we denote the local empirical loss function of client $i$ by:
@@
$\mathbb{E}\left[\frac{1}{|\mathcal{B}|}\sum_{(\mathbf{x},y)\in\mathcal{B}}\nabla_{\mathbf{w}}l(h((\mathbf{x};\mathbf{w}),y))\right]=\nabla J_{i}(\mathbf{w}),\forall i\in\mathcal{N}.$ (23)
@@
$\mathbb{E}\left[\left\|\frac{1}{|\mathcal{B}|}\sum_{(\mathbf{x},y)\in\mathcal{B}}\nabla_{\mathbf{w}}l(h((\mathbf{x};\mathbf{w}),y))-\nabla J_{i}(\mathbf{w})\right\|_{2}^{2}\right]\leq\sigma^{2},\forall i\in\mathcal{N}.$ (24)
The data distributions on clients vary among each other, as explained in Section 3.1, which leads to different optimization objectives and diverse gradients. To show such data heterogeneity, previous works [47, 48] made assumptions on the difference between local and global training objectives. Different from these works, we aim to investigate the collaboration among clients and thus make a more precise assumption on the gradient similarity among clients as follows.
@@
We are now to derive the convergence of the HCCT in the following theorem. Previous works in FL [49, 5, 50] consider that an algorithm has reached convergence if it converges to a stationary point of the global loss function, i.e., if its expected squared gradient norm $\min_{t\in[T]}\mathbb{E}[\|\nabla J_{i}(\mathbf{w})\|^{2}]$ is zero [51]. Compared with these works, the clients in cross-silo FL have local training objectives and are not concerned about the global objective. Therefore, a better option is to verify the convergence of the algorithm by upper bounding the sum of squared norm of local gradients $\sum_{i\in\mathcal{N}}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})\|^{2}]$ [52].
@@
Let $\xi_{T}=\sum_{t=0}^{T-1}\eta^{t}$. Consider client $i$ joins in group $G_{i}^{t}$ at epoch $t$. With Assumptions 1-3, if the learning rates satisfy $\eta^{t}<\frac{1}{L},\forall t\in[T]$, we have:
@@
$+\frac{1}{\xi_{T}}\sum_{t=0}^{T-1}2\eta^{t}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$
$+\frac{1}{\xi_{T}}\sum_{t=0}^{T-1}2\eta^{t}L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{i\in\mathcal{N}}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}$
$+\frac{1}{\xi_{T}}\sum_{t=0}^{T-1}L(\eta^{t})^{2}N\sigma^{2}+\sum_{t=0}^{T-1}4\eta^{t}L^{2}N\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2}.$ (26)
@@
We first prove an upper bound for the local loss decay of all clients at training epoch $t$ (formally stated in Lemma 1 in Appendix B), which is given by:
$\leq$ $-\eta^{t}\sum_{i\in\mathcal{N}}\mathbb{E}\langle\nabla J_{i}(\mathbf{w}_{i}^{t}),\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\rangle$
$+\frac{L(\eta^{t})^{2}}{2}\sum_{i\in\mathcal{N}}\mathbb{E}\left[\left\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\right\|^{2}\right]+\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}.$ (27)
In the right-hand side (RHS) of (27), the inner product $-\eta^{t}\mathbb{E}\langle\nabla J_{i}(\mathbf{w}_{i}^{t}),\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\rangle$ needs to be further upper bounded. The main challenge is to bound the divergence between any two local gradients, i.e., $\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$. Using $L$-smoothness in Assumption 1, we show that:
@@
Then we observe that, despite evolving from different groups, $\mathbf{w}_{j}^{t}$ and $\mathbf{w}_{i}^{t}$ have the same initialization $\mathbf{w}^{0}$. Therefore, we can upper bound their difference by comparing the accumulated gradients in previous $t-1$ iterations.
Finally, the result is completed by summing up both sides of an upper bound of (27) over index $t=0,1,\dots,T-1$ and dividing them by $\sum_{t=0}^{T-1}\frac{\eta^{t}}{2}$. The detailed proof is deferred to Appendix C. ∎
@@
When the learning rates satisfy $\lim_{T\rightarrow\infty}\sum_{t=0}^{T-1}\eta^{t}=\infty$, $\lim_{T\rightarrow\infty}\sum_{t=0}^{T-1}(\eta^{t})^{2}<\infty$, and $\eta^{t}=\mathcal{O}\left((\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2})^{1/p}\right)$ with $p>0$, the RHS of (26) converges to zero as $T\rightarrow\infty$, i.e., the local models output by HCCT converge to the stationary points of the local loss functions.
Proof.
Please refer to Appendix D. ∎
From Theorem 2 and Corollary 1, we observe that the convergence of HCCT is hindered by the gradient dissimilarity among clients. Specifically, as more heterogeneous clients are grouped, i.e., $\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$ is larger, the convergence speed will be slowed down. This verifies our intuition that clients with similar gradients should be divided into the same collaboration group. In the next section, we will show the empirical benefits of the proposed HCCT scheme via extensive simulations.
Note that we adopt a fixed learning rate during a local epoch and time-varying learning rates across (global) training epochs. This setup follows previous studies on FL e.g., [1, 47], and studies on local SGD, e.g., [53, 54]. As we focus on comparing the impact of data distributions and dataset sizes, we adopt the simple but classic SGD optimizer as the client optimizer and leave the investigation of HCCT with varying local learning rates as future work.
@@ 7.1 Setup
  • CIFAR-10 [34]: In this training task, $N=10$ clients have heterogeneous and imbalanced data samples. We divide the training data into 100 shards, each of which contains one random class of samples. Then, we randomly allocate a certain number of shards to each client using a half-normal distribution $\mathcal{H}(1)$.
@@
  • FedFA [^2] This work [25] adopts SGD with momentum as local training algorithm, but for fair comparisons we use mini-batch SGD in all methods. [25]: All clients train in a global model similar to the global training scheme. The key difference is that their model updates are weighted according to the training accuracy ${A}_{i}^{t}$ on local data, i.e., $\hat{\mathbf{g}}_{1}^{t}=\sum_{i\in\mathcal{N}}\frac{-\log_{2}{A}_{i}^{t}}{\sum_{j\in\mathcal{N}}-\log_{2}{A}_{j}^{t}}\mathbf{g}_{i}^{t}$, which ensures equitable accuracy distribution among clients.
@@
IFCA and FLSC serve as benchmarks for client clustering in FL, while FedFA and MAXFL represent other approaches to determining client collaboration patterns. To demonstrate the compatibility of HCCT, we also show the simulation results for two variants: HCCT-E (an efficient implementation in Section 5.2) and HCCT-P (with personalization layers [23]).
@@ 7.2 Performance Comparison
In the FMNIST training task, HCCT and global training have similar test errors, and they outperform other baselines. Given the IID data at clients, a global collaboration incorporates the maximum training data samples and significantly improves the generalization performance of the model. However, clients in the CIFAR-10 task have heterogeneous data, and thus the global training scheme suffers from such non-IID issues. In this case, clients should perform independent training, and HCCT also exhibits comparable test error to independent training. These results evidence that HCCT can identify the most effective client collaboration pattern, including independent training and global training, to accommodate various scenarios.
To further compare different training schemes, we summarize the average, standard deviation, minimal, and maximal values of test errors at clients in the Digit training task in Table V. We see that in addition to achieving the lowest average test error, HCCT also notably reduces the test error of the worst-performing client, which verifies that clients with fewer training data samples can benefit from effective collaboration. Meanwhile, the generalization performances of those clients with sufficient training data are also preserved, as the proposed utility function excludes clients with dissimilar data from its collaboration group. Besides, it is worth noting that HCCT achieves a lower standard deviation than the baseline schemes, which shows that HCCT implicitly enhances fairness among clients, leading to similar generalization performance for all clients.
@@ 7.4 Effect of System Parameters
Effect of $\alpha$. In Fig. 4, we investigate the effect of $\alpha$ using the CIFAR-10 dataset. A larger value of $\alpha$ emphasizes the importance of data volume, while a small value means that clients tend to collaborate with someone with similar data distribution. In this study, we split the assigned data for each client into the training data and test data according to different ratios, where a larger ratio means that a client has more training data samples but less test data. In general, when $\alpha$ is small, its effect on the model performance is rather trivial. We observe from the results that as the ratio becomes larger, more training data speed up the training process and enhance the generalization performance in all cases. Moreover, if clients have fewer training data samples (i.e., the ratio is $0.05$ or $0.1$), the optimal value of $\alpha$ is $100$. It means that clients have limited local data and tend to collaborate with others to benefit from their training data. In contrast, when clients have many local data samples (i.e., ratio is $0.2$ or $0.5$), they will emphasize more on the data distribution. Thus, by using a smaller $\alpha$ clients can avoid collaborating with other clients whose data distribution diverges significantly. These results provide a guideline for the selection of the value of $\alpha$ in the utility function. In addition, it is straightforward to extend and apply personalized $\alpha$ for different clients.
@@
Effect of the number of groups. Next, to study how the number of groups affects the generalization performance, we set the number of clients as $N=20$, manually fix the number of groups as $K$, and evaluate the test error in the Digit training task. From the results in Fig. 5, we observe that in this scenario the optimal number of groups to achieve the minimal test error is $10$. Notably, the maximal utility is also achieved when clients are divided into $10$ groups. A plausible explanation is that each group consists of clients sharing the same dataset. This implies that the utility function can effectively reflect the client’s requirement of improving the generalization performance, such that the minimal test error is achieved when the utility is maximized.
Furthermore, we set the number of clients as $N\in\{10,20,30,40,50\}$, manually fix the number of groups as $K\in\{1,2,5,10,N\}$, and evaluate the test error in the Digit training task. The results in Fig. 6 show that the clients prefer collaborating with a limited number of clients instead of joining in a global federation.
Figure 6: Local test error (mean) in the Digit dataset with different numbers of groups $K$ and clusters $N$.
@@ 8 Conclusions
In this work, we tackled the challenge of optimizing the client collaboration pattern in order to maximize the generalization performance in cross-silo FL. We derived the generalization bound for clients in various collaboration cases and then formulated the client utility maximization problem. To efficiently solve this problem, we proposed HCCT, a hierarchical clustering-based collaborative training scheme, in which clients are partitioned into different non-overlapping groups without the need to initially decide the number of groups. We also proved the convergence of HCCT for general loss functions. The effectiveness of HCCT is further verified via extensive simulations in different scenarios and datasets.
The group partition function in HCCT requires evaluating client similarity for computing the client utility. This introduces additional computational costs on the central server, increasing with the number of clients. Given this scalability issue, HCCT is limited to the cross-silo FL with fewer clients. For future work, it is interesting to design the collaboration patterns for cross-device FL [4, 61]. Besides, it is worth exploring personalization techniques in HCCT to further improve the generalization performance. In addition, adapting HCCT to the case with time-varying local learning rates would be beneficial.